# S2CFormer: Reorienting Learned Image Compression from Spatial Interaction to Channel Aggregation

### v2

@@ 4.2 Rate-Distortion Performance
The performances of our models suggest that complex spatial interaction design may be redundant and can be simplified. S2C-Conv and S2C-Attention show near-identical performance, indicating a specific choice of spatial interaction is not critical for LIC. Instead, FFN-based channel aggregation is the key factor that secures R-D performance.
@@
Ablations of Channel Aggregation.  In this section, we further analyze the critical role of Channel Aggregation within our S2CFormer architecture. To assess their impact, we first conducted an ablation study by removing all FFN modules from S2C-Conv and S2C-Attention.
Fig. 4 illustrates the R-D curves for this ablation study. Notably, pure separable convolution and window attention are significantly less effective than their FFN-enhanced counterparts. Specifically, adding FFNs to separable convolution improves BD-rate by -10.57%, while FFNs in S2C-Attention lead to an -11.14% improvement. These results highlight the substantial contribution of FFNs to the overall R-D performance of our LIC models.
In contrast, S2C-Identity, which eliminates spatial interaction while retaining channel aggregation, approaches the performance of TCM-large, outperforming the two models without CA by -7.51% and -8.17% BD-rate respectively. This comparison underscores the greater importance of FFNs in LIC models.
This ablation study strongly supports using FFNs for channel aggregation in future LIC models.
@@
Figure 10: Visual Comparison. This figure presents visualizations of decompressed images from the Kodak dataset using various methods. Each subfigure is labeled with “Method |||| Bpp |||| PSNR |||| MS-SSIM”.
We also examine the ERF [36] of analysis networks to demonstrate the significance of channel aggregation. Although it is commonly believed that the attention mechanism expands the receptive field and aggregates more non-local information, our findings challenge this view. As shown in Fig. 5, both the S2C-Conv and S2C-Attention models, when lacking channel aggregation, exhibit smaller ERF. In contrast, when channel aggregation is introduced, the ERF of both models expands to cover more areas of the input image. We also evaluate the Swin Transformer-based LIC model and observe that window shifting does not significantly enlarge the ERF. Similar to the S2C-based models, Swin-based models also rely on channel aggregation to broaden ERF. This demonstrates that channel aggregation is the key factor in activating more pixels in the original image and capturing additional information, which boosts the R-D performance of LIC models. These comparisons suggest that some complex and time-consuming spatial operations appear redundant.
@@
Performance of S2C-Hybrid.  We enhance model performance by integrating the advantages of different S2C-instantiations and advanced FFN. We increase $L_{2}$ and $L_{3}$ to 8 while keeping other configurations unchanged. To manage decoding latency and complexity, we utilize the S2C-Conv block in the initial stage to reduce complexity; in subsequent stages, we employ the S2C-Attention block. As shown in Fig. 6-8, by employing only the fundamental S2CFormer blocks with simplified spatial interaction modules, our S2C-Hybrid model surpasses all other LIC models in R-D performance, exceeding VTM-21.0 by 14.28%, 12.88%, and 17.20% in BD-rate for the Kodak, CLIC, and Tecnick datasets, respectively. This result demonstrates the effectiveness of integrating various simple spatial interactions within a unified architecture to optimize both computational efficiency and model performance.
Ablations on S2C-Hybrid For the three-stage main transform networks, we employ S2C-Conv for the first stage and S2C-Attention for the others, which is the best arrangement according to our experiments. As shown in Tab.2, we represent S2C-Conv and -Attention with “C” and “A” respectively. “[C,A,A]” appears to be the best, the reason for which may be that convolution can better capture local details at high resolution, while attention excels at global information when resolution drops.
@@
We also have experimented about the S2C-Hybrid model’s complexity in Tab.3. The base model Hybrid-S has 3 blocks for each stage. By stacking more blocks in the last two stages (adding blocks to the first stage greatly slows decoding), performance steadily improves.
Stages[1,2,3] Hybrid-S[3,3,3] Hybrid-M[3,5,5] Hybrid-L[3,8,8]
BD-rate (%) -13.35 -13.85 -14.28
FLOPs (T) 3.20 3.59 4.17
Decoding (s) $\downarrow$ 0.357 0.358 0.362
Table 3: S2C-Hybrid Complexity Ablations
Visual Comparison.  We present the visualization results in Fig. 10, comparing our method with the traditional codec VTM 21.0 and the advanced LIC model TCM.
@@ 4.3 Model efficiency
We calculate the FLOPs and decoding latency for various models using 2K images, and evaluate training throughput with a batch size of 8 and patch size of 256×256. Training throughput is directly correlated to the training time of LIC models. Numeric results are provided in Table 1.
@@
FLOPs.  For VAE-based LIC models, FLOPs are significantly influenced by the number of channels in the first stage. We directly set all channel numbers as 192, following the original setting of LIC model [2], as the optimal channel settings are beyond the scope of this study. Some leading methods have reduced the first-stage channel number to 128 or fewer [43, 27], which significantly lowers FLOPs. However, these models are still much slower than our S2CFormer-based models in both training and decoding speed, which are the primary concerns for LIC.
5 Conclusion and Future Work
We revisit the key factors that enhance the R-D performance of transformer-based LIC models. Specifically, we identify certain spatial operations in previous leading methods as redundant. To address this, we introduce S2CFormer, a novel paradigm for nonlinear transform blocks in LIC. This architecture consists of two core components: the Spatial Interaction module and the Channel Aggregation module. By shifting the focus of LIC from spatial interaction to channel aggregation, this approach aims to optimize performance while also improving speed. Through extensive experiments and comparisons, we demonstrate that FFN-based channel aggregation is critical for improving LIC performance. Models based on S2CFormer offer faster training speed and lower decoding latencies compared to existing methods while maintaining competitive R-D performance. Moreover, we suggest that FFN-based channel aggregation remains a promising avenue for future research.
By combining simple spatial interaction operators with the proposed Gated-FFN for channel aggregation, our models achieve state-of-the-art R-D performance on the Kodak, Tecnick, and CLIC Professional Validation datasets, while also maintaining efficiency. This approach breaks the suboptimal trade-off between R-D performance and decoding latency inherent in previous methods.

### v3

@@ 4.2 Basic S2CFormer Instances
The performances of our models suggest that complex spatial interaction design may be redundant and can be simplified. S2C-Conv and S2C-Attention show near-identical performance, indicating a specific choice of spatial interaction is not that critical for LIC. Instead, FFN-based channel aggregation is the key factor that secures R-D performance.
@@
4.3 Channel Aggregations
Ablations of Channel Aggregation.  In this section, we analyze the critical role of Channel Aggregation (CA) in the S2CFormer architecture. An ablation study was conducted by removing all FFN modules from S2C-Conv and S2C-Attention. Results in Fig. 5 show that FFN-enhanced separable convolution and window attention significantly outperform their counterparts, improving BD-rate by -10.57% and -11.14%, respectively. This highlights the substantial impact of FFNs on R-D performance. In contrast, S2C-Identity, which retains CA but eliminates spatial interaction, approaches TCM-large’s performance, outperforming models without CA by -7.51% and -8.17% BD-rate. The study emphasizes the importance of FFNs in LIC models and supports their use for CA in future designs.
@@ 4.3 Channel Aggregations
Figure 11: Visual Comparison. This figure presents visualizations of decompressed images from the Kodak dataset using various methods. Each subfigure is labeled with “Method |||| Bpp |||| PSNR |||| MS-SSIM".
We analyze the Effective Receptive Field (ERF) [37] of networks to highlight the importance of channel aggregation. Contrary to the belief that attention mechanisms expand the receptive field and aggregate non-local information, our findings show that both S2C-Conv and S2C-Attention models exhibit smaller ERFs without channel aggregation. As shown in Fig. 6, introducing channel aggregation significantly expands the ERF, covering more input image areas. Similarly, Swin Transformer-based LIC models rely on channel aggregation to broaden ERF, as window shifting alone does not significantly enlarge it. These results indicate that channel aggregation is crucial for activating more pixels and enhancing R-D performance, suggesting that certain complex spatial operations may be redundant.
@@
4.4 S2C-Hybrid
Arrangement Ablations. Our experiment results in Tab. 2 show that the optimal configuration for the three-stage main transform networks is to use S2C-Conv (denoted as “C”) in the first stage and S2C-Attention (denoted as “A”) in the subsequent stages. The “[C, A, A]” configuration performs best, likely because convolution layers excel at capturing local details at high resolutions, while attention mechanisms are more effective at extracting global information as resolution decreases. This highlights that while S2C-Conv and S2C-Attention exhibit similar rate-distortion performance, their distinct capabilities at different stages significantly impact overall performance when optimally arranged.
@@ 4.4 S2C-Hybrid
Performance and Scaling-Up. As shown in Table 1, the standard S2C-Hybrid model variant, Hybrid-S, which uses a configuration of {3, 3, 3} for stages $L_{1}$, $L_{2}$, and $L_{3}$, has achieved state-of-the-art performance across all three datasets while maintaining low decoding latency. Subsequently, we scaled up the S2C-Hybrid model to further explore the trade-off between model complexity and rate-distortion performance. Hybrid-M uses {3, 5, 5} blocks and Hybrid-L uses {3, 8, 8} blocks for stages $L_{1}$, $L_{2}$, and $L_{3}$. As reported in Table 1, stacking more blocks in the last two stages consistently improves performance. The RD curves for Hybrid-L are presented in Figures 79.
Visual Comparison As illustrated in Fig. 11, comparisons with VTM-21.0 and TCM demonstrate that our proposed method achieves superior detail preservation.
@@ 4.5 Model efficiency
We calculate the FLOPs and decoding latency for various models using 2K images, and evaluate training throughput with a batch size of 8 and patch size of 256$\times$256. Training throughput is directly correlated to the training time of LIC models. Numeric results are provided in Table 1.
@@
FLOPs.  For VAE-based LIC models, the number of channels in the first stage significantly influences the FLOPs. In our work, we set all channel numbers to 192, following the classical setting of LIC models [2, 42, 3, 10, 21], since optimizing channel settings falls beyond the scope of this study. Although some leading methods reduce the first-stage channel count to 128 or fewer [44, 28] to lower FLOPs, these models remain considerably slower than our S2CFormer-based models in both training and decoding.
This paper focuses on rebalancing spatial interactions and channel aggregation for low decoding latency. In this context, S2CFormer serves as a paradigm that permits flexible channel adjustments to reduce FLOPs. To demonstrate this, we propose a variant named Hybrid-T, designed for low-FLOP operation. Specifically, we set the channel numbers $C_{1}$, $C_{2}$, and $C_{3}$ to 96, 192, and 256, and the block numbers $L_{1}$, $L_{2}$, and $L_{3}$ to 3, 5, and 8, respectively. Hybrid-T achieves a SOTA BD-rate of -13.32% on the Kodak dataset, while its FLOPs are only 2.63T—a value relatively low compared to previous SOTA methods. These results indicate that our S2CFormer is not a fixed design, and we can achieve even more efficient S2C models by adjusting the channel numbers to find the optimal configuration in the future.
5 Conclusion
This paper tackles the trade-offs between decoding latency and rate-distortion performance in transformer-based LIC models. We emphasize the critical role of channel aggregation and demonstrate through extensive experiments that channel operations—rather than some time-consuming spatial interactions—are key to achieving high-performance LIC models. Based on these insights, we propose the S2CFormer paradigm, which streamlines spatial operations while substantially enhancing channel aggregation.
Building on these findings, we evaluate several S2CFormer instances. Experimental results from S2C-Conv and S2C-Attention confirm that combining simpler spatial interactions with effective channel aggregation not only delivers state-of-the-art compression performance but also significantly accelerates decoding speed. Furthermore, the advanced S2C-Hybrid model, which integrates the strengths of multiple S2C variants, establishes a new benchmark for LIC by outperforming existing models on standard datasets such as Kodak, Tecnick, and CLIC Professional Validation. Overall, this work sets a new standard for efficient, high-performance LIC and lays a robust foundation for future research on optimized channel aggregation strategies.